OpenStudy (anonymous):

find the square root of 5184

OpenStudy (phi):

without using a calculator? if so, find the prime factors. Start by dividing by 2 as many times as you can. 5184= 2*2592 now divide 2592 by 2 we get 5184= 2*2*1296 keep going. then try 3, 5, 7, 13 (the primes) until you end up with all prime factors. can you do that ?

OpenStudy (mrnood):

If no calculator then you could use logs - but you probably need calculator for that (there used to be books with logs in believe it or not!) If by pencil and paper then use @phi 's method

OpenStudy (anonymous):

You can also approximate at first, thinking of square numbers close to 5184, such as \(4900=70^2\), then keep moving up the number line.

OpenStudy (anonymous):

The number in the one's digit should also give you an indication of what to look for. Which numbers when squared give a 4 in the one's place? 2 and 8 work, so I would guess either 72 or 78.